OpenStudy (anonymous):

Need to solve graphically by first converting to an equivalent equation with 0 on the right-hand side and then find the x-intercepts. |2x-5|=4-|x-3| HELP!?

OpenStudy (anonymous):

There are 4 casses

OpenStudy (anonymous):

\[\begin{split} -(2x-5)&=4-(x-3)&\quad&2x-5<0&\quad&x-3>0\\ -(2x-5)&=4+(x-3)&&2x-5<0&&x-3<0\\ 2x-5&=4-(x-3)&&2x-5>0&&x-3>0\\ 2x-5&=4+(x-3)&&2x-5>0&&x-3<0 \end{split} \]So there are \(4\) possible solutions. You have to find them all.

OpenStudy (anonymous):

Some of these cases might actually be impossible though.